OpenStudy (anonymous):

What is the equation of the line with the coordinates (8,0) (0,8)?

OpenStudy (anonymous):

y=-1x+c

OpenStudy (kira_yamato):

slope = (8-0)/(0-8) = -1 Let y intercept = c 0 = -8 + c c = 8 Eq'n: y = 8-x

OpenStudy (anonymous):

thanks! (: